OpenStudy (henrietepurina):

The tables below show the values of f(x) and g(x) for different values of x:

OpenStudy (henrietepurina):

f(x) = 2(3)x x f(x) -2 0.22 -1 0.67 0 2 1 6 2 18

OpenStudy (henrietepurina):

g(x) = 3x + 9 x g(x) -2 9.11 -1 9.33 0 10 1 12 2 18

OpenStudy (henrietepurina):

Based on the tables, what is the solution to the equation 2(3)x = 3x + 9?

OpenStudy (henrietepurina):

x = 0 x = 2 x = 12 x = 18

OpenStudy (henrietepurina):

@undeadknight26 can you help or you dont know this?

undeadknight26 (undeadknight26):

I really have no idea what this question is asking...I get x=3 though...

OpenStudy (henrietepurina):

neither do i :(

undeadknight26 (undeadknight26):

@amistre64 @aaronq @surjithayer @e.mccormick @razor99 @whpalmer4 @iPwnBunnies @ParthKohli

Parth (parthkohli):

Basically, the question is asking you the values of \(x\) for which \(f(x) = g(x)\).

Parth (parthkohli):

Look at the table. For which value of \(x\) in the table are \(f(x)\) and \(g(x)\) the same?

OpenStudy (henrietepurina):

2 18?

Parth (parthkohli):

That's right! So the solution is \(x =2\).

OpenStudy (henrietepurina):

oh, thanks :)

OpenStudy (anonymous):

i think problem is f(x)=g(x) based on the tables \[f(x)=2\left( 3 \right)x=2\left( 3 \right)^x\] \[g \left( x \right)=3x+9=3^x+9\] f(x)=g(x) gives \[2*3^x=3^x+9\] \[2*3^x-3^x=9\] \[\left( 2-1 \right)3^x=9=3^2,3^x=3^2,x=2\]